OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge y = x^2 \sin x \tan x\] product rule??

OpenStudy (lgbasallote):

do i change it to \[\huge y = x^2 (\frac{\sin^2 x}{\cos x})\] then use product rule?

OpenStudy (anonymous):

Ya i would try then lgba

OpenStudy (anonymous):

So we will let f (x) = x^2sin(x)tan(x) so the derivative will require the product rule!

OpenStudy (anonymous):

But usually when we do a product rule we are only working with 2 expressions. Here we have three though. So this rule looks like for\[ \huge\ f(x) = a*b*c \]where a, b, and c are all functions of x, we have: \[\huge\ f '(x) ab*d(c)/dx + a*d(b)/dx*c + d(a)/dx*bc\]

OpenStudy (anonymous):

I am going correctt @lgbasallote

OpenStudy (lgbasallote):

so how to product rule this...

mathslover (mathslover):

i think u r going right @Rohangrr

OpenStudy (lgbasallote):

what is the derivative of \[\frac{\sin^2 x}{\cos x}\]

OpenStudy (anonymous):

So we have\[ f '(x) = x^2\sin(x)*d(tanx)/dx + x^2*d(sinx)/dx*\tan(x) + d(x^2)/dx*\sin(x)\tan(x)\] So \[f '(x) = x^2\sin(x)\sec^2(x) + x^2\cos(x)\tan(x) + 2xsin(x)\tan(x) \] So \[f '(x) = x^2\sin(x)\sec^2(x) + x^2\sin(x) + 2xsin(x)\tan(x) \] final answer \[\huge\ =\sin(x)[x^2\sec^2(x) + x^2 + 2xtan(x)]\]

OpenStudy (lgbasallote):

lol wow nice

mathslover (mathslover):

it looks great @Rohangrr

OpenStudy (anonymous):

medals

OpenStudy (lgbasallote):

let's see if it's right...

OpenStudy (anonymous):

lol

mathslover (mathslover):

if some one blocks then how to give medal ?

OpenStudy (anonymous):

undertaker

OpenStudy (anonymous):

there is a easier mwthod deffrentiate both wid respect to dx

OpenStudy (lgbasallote):

\[\huge \frac{x^2}{\sin x} + x^2 \sin x + \frac{2x \sin^2 x}{\cos x}\]

OpenStudy (lgbasallote):

lol wow...this is a hard integral...

OpenStudy (anonymous):

k so lgba i am correct

OpenStudy (lgbasallote):

like i said..im going to check first

mathslover (mathslover):

yes u r right @lgbasallote first check :D

OpenStudy (lgbasallote):

heh this is hard interal...im just gonna derive..

mathslover (mathslover):

best of luck my friend :)

OpenStudy (anonymous):

thanks saif u forgot bout the mail

OpenStudy (saifoo.khan):

@Rohangrr , OHHHH! Man! My bad. I totally forgot. Please forgive me! :'( I will do it soon! :D

OpenStudy (lgbasallote):

yup @Rohangrr made a mistake

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

np

OpenStudy (saifoo.khan):

^_^

OpenStudy (lgbasallote):

pls check guys...here's what i got \[\huge \sin x (2x^2 + 3x\tan^2 x)\] am i right or @Rohangrr ?

OpenStudy (anonymous):

what do u mean @LGBA "am i right or @Rohangrr""

OpenStudy (lgbasallote):

we have different answers...so who's right

OpenStudy (lgbasallote):

lol @saifoo.khan

OpenStudy (lgbasallote):

idk the answer too im just checking my work

OpenStudy (saifoo.khan):

;)

OpenStudy (anonymous):

saif is ........ hmmmm........ why were u worrying so mch @lgbasallote

OpenStudy (lgbasallote):

worrying?

OpenStudy (anonymous):

chillllll okay okay leave it

OpenStudy (lgbasallote):

lol so which is correct?

OpenStudy (lgbasallote):

im trying to see if they're actually equal...

OpenStudy (lgbasallote):

oops i made a mistake...but it's still not similar to @Rohangrr 's answer... \[\huge \sin x (2x^2 + \tan^2 x + 2x\tan x)\]

OpenStudy (anonymous):

wait in that some part is hiding i ll do it

OpenStudy (mimi_x3):

waitttt why not change i to.. \[y = x^2sinxtanx => x^2*\frac{1-\cos^{2}x}{cosx} => x^2*\frac{1}{cosx} - \frac{\cos^{2}x}{cosx} => x^2secx-cosx\]

OpenStudy (mimi_x3):

\[y = x^2sinxtanx => x^2*\frac{1-\cos^{2}x}{cosx} => x^2*\frac{1}{cosx} - \frac{\cos^{2}x}{cosx} =\] \[ => x^2secx-cosx\]

OpenStudy (anonymous):

my 3rd conversation where a, b, and c are all functions of x, we have: \[f '(x) ab*d(c)/dx + a*d(b)/dx*c + d(a)/dx*bc \]

OpenStudy (mimi_x3):

what is that? lol

OpenStudy (mimi_x3):

btw, i want to say something off topic; am i the only who is experiencing problems with this site???

OpenStudy (lgbasallote):

nope

OpenStudy (lgbasallote):

so now we have 3 different answ

OpenStudy (anonymous):

well other way is: \[\huge\ ycos(x) = x^2\sin^2(x) \] Using implicit differentiation treating y as a function of x we have: \[y(-sinx) + \cos(x)dy/dx = x^2[2\sin(x)\cos(x)] + 2xsin^2(x) \] So isolating dy/dx on the left side of the equation: \[\cos(x)dy/dx = x^2\sin(2x) + 2xsin^2(x) \] So \[\huge dy/dx = [x^2\sin(2x) + 2xsin^2(x)] / \cos(x)\]

OpenStudy (mimi_x3):

noooo..im suggesting you to change it to \[ x^2secx-cosx\] and apply the product rule; it wont look messy why implicit differentiation????

OpenStudy (lgbasallote):

yup it turnsout as a completely new answer4

OpenStudy (lgbasallote):

that 4 does not mean anything

OpenStudy (anonymous):

lol u all are getting confused and confusing me

OpenStudy (lgbasallote):

so who's right??

OpenStudy (anonymous):

I think question by answer not medal

OpenStudy (lgbasallote):

lol what?

OpenStudy (mimi_x3):

wait. is this for a bigger problem?

OpenStudy (lgbasallote):

i believe so..

OpenStudy (anonymous):

QUESTION --> answer QUESTION -\-> medal

OpenStudy (lgbasallote):

i still dont get what you mean

OpenStudy (lgbasallote):

hmm i giuess no one can do it haha

OpenStudy (mimi_x3):

what is the problem? my method would work; however, tell us what's the problem that you're working on?

OpenStudy (lgbasallote):

lol look at the blue box :p

OpenStudy (mimi_x3):

lol i know..the bigger problem towards the question? that question must have been derived from something

OpenStudy (lgbasallote):

the book :P lol